These were accepted.
Alison has circulated around CEG and has incorporated comments. She'll now publish to COs. Also merge "procedures" action below.
These are on-line somewhere now. Alison will check that they're linked in an easy-to-find way.
Security have agreed. Dave H apparently has a key. Alastair to check with him.
Merge this action with "publish checklist for firewall holes".
Toby is arranging a meeting. Amend action accordingly.
Alison, Graham, Lindsey, Neil and Toby met. Notes are here.
All Units should now review and act on the results.
New scripts done. Iain now to find a good home for them, and to write some documentation.
Ian to bundle this into DevProj#279
Blog articles suggested:
Blog articles completed:
Report from Computing Executive Group
Reports from units
Dave R has escalated the Server Room UPS situation to David Barratt. Now waiting...
rattle: this looked as though something was holding a kernel lock somewhere. Perhaps there's a fix for something in a later kernel version?
blatiere: the "context change timeout" message was essentially an indication that lcfg-client's test that the daemon had started properly had failed. It didn't appear to be conserver-related, as that's started later in the boot sequence. Ian and Stephen will keep an eye on it.
The new script will be on all DICE machines, from next week's <stable> release.
Topics for discussion
Craig exhorted everyone to add suggestions for next week's technical disussion meeting to the wiki page.
The next meeting will be on 29th January 2014 at 10:00 in IF-1.15.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh